OpenStudy (anonymous):

Help me!!!! the dimensions of a rectangle are such that its length is 5 in. more than its width. If the length were doubled and if the width were decreased by 2 in., the area would be increased by 162 in squared. What are the length and width of the rectangle?

OpenStudy (anonymous):

I got W = 14 in. and L = 19 in. I could be wrong...

OpenStudy (anonymous):

I haven't gotten anything because this is the equation I came up with (2(x+5))(x-2)=162 and now can't solve it lol

OpenStudy (anonymous):

You actually have three equations. The equation you got would mean that the area is 162 instead of being INCREASED by 162. Equation 1: You know that the length is 5 more than the width. \[L=W+5\] Equation 2: Equation for original area: \[A=LW\] Equation 3: The length doubled (2L) times the width decreased by 2 (W-2) would give you 162 more than the original area (A+162) \[(2L)(W-2)=A+162\] So here are the 3 equations: \[L=W+5\]\[A=LW\]\[(2L)(W-2)=A+162\] Substitute Equation 2 into Equation 3: \[(2L)(W-2)=LW+162\]Substitute W+5 for L: \[2(W+5)(W-2)=W(W+5)+162\]Expand both sides: \[2W^2+6W-20=W^2+5W+162\] Put this into the form 0=ax^2+bx+c \[0=W^2+W+182\] You can factor this: \[0=(W+14)(W-13)\] This gives the two solutions W=13 and W=-14. W=-14 doesn't make sense, because the width must be positive. Therefore, your width is 13. Add 5 to get the length (18).